# arXiv:2603.03220v1[cond-mat.mtrl-sci] 03 Mar 2026

### v1

@@ Appendix A Källén-Lehmann spectral representation and the
The numerical evaluation of the density response functions is a notoriously difficult undertaking and the following relations [Nyquist1928, Callen1951, Fetter1971, Kubo1966, Abrikosov1975, Giuliani2005] appeared useful in the validation of our computer implementation and the interpretation of results. The relations given below are completely general and apply even to the cases of NCity and non-centrosymmetric crystals with relativistic effects taken into account.
@@
which in turn implies that $\chi_{\alpha\beta}^{\mathcal{R}}{\left({\mathbf{q}},\omega=0\right)}$ is a Hermitian and $\chi_{\alpha\beta}^{\mathcal{R}}{\left({\mathbf{q}}={\mathbf{0}},\omega=0\right)}$ a symmetric matrix.

### v2

@@ Appendix A Källén-Lehmann spectral representation and the
The numerical evaluation of the density response functions is a notoriously difficult undertaking and the following relations [Nyquist1928, Callen1951, Fetter1971, Kubo1966, Abrikosov1975, Giuliani2005] appeared useful in the validation of our computer implementation and the interpretation of results. The relations given below are completely general and apply even to the cases of non-collinearity non-centrosymmetric crystals with relativistic effects taken into account.
@@
which in turn implies that $\chi_{\alpha\beta}^{\mathcal{R}}{\left({\mathbf{q}},\omega=0\right)}$ is a Hermitian and $\chi_{\alpha\beta}^{\mathcal{R}}{\left({\mathbf{q}}={\mathbf{0}},\omega=0\right)}$ a symmetric real matrix.